I'm having killer difficulty figuring out Solving Equations with More than One Operation....
Here is an example:
8x+3-x=5+4x-2

Answer by mananth(16946)

I'm having killer difficulty figuring out Solving Equations with More than One Operation....
Here is an example:
8x+3-x=5+4x-2
Generally the method of solving such eauations is to bring the terms on the left hand side of the equation and the numbers on the right sude of the equation.
Rule 1 The sign of the number or term changes if it moves to the other side of the equation.
Rule 2. the sign changing is applicable only for - and +.
8x-x= 5+4x-2 -3. I have taken 3 from left to right. so it changes sign
8x-x-4x=5-2-3. 4x goes from right to left . sign has changed
3x=0
x=0/3
x=0